OpenStudy (anonymous):

What is the torque generated at the elbow by a 90-N force pulling on the forearm at an angle of 110deg from the horizon at a point 18cm from the elbows axis of rotation. The forearm is postioned horizontal to the ground

OpenStudy (anonymous):

Take the cross product of the radius and the force. \[\tau = r \times F\]Since we are in 2-D, this becomes\[\tau = F r \sin(\theta)\] where theta is the angle between the arm and the force (110 degrees), F is the force, and r is the moment arm (0.18 m)